# Reffly: Melody-Constrained Lyrics Editing Model

### v1

@@ Abstract
Automatic melody-to-lyric generation aims to produce lyrics that align with a given melody. Although previous work can generate lyrics based on high-level control signals, such as keywords or genre, they often struggle with three challenges: (1) lack of controllability, as prior works are only able to produce lyrics from scratch, with little or no control over the content; (2) inability to generate fully structured songs with the desired format; and (3) failure to align prominent words in the lyrics with prominent notes in the melody, resulting in poor lyrics-melody alignment. In this work, we introduce Reffly (REvision Framework For Lyrics), the first revision framework designed to edit arbitrary forms of plain text draft into high-quality, full-fledged song lyrics. Our approach ensures that the generated lyrics retain the original meaning of the draft, align with the melody, and adhere to the desired song structures. We demonstrate that Reffly performs well in diverse task settings, such as lyrics revision and song translation. Experimental results show that our model outperforms strong baselines, such as Lyra Tian et al. (2023) and GPT-4, by 25% in both musicality and text quality.
@@ 1 Introduction
Automatic melody-to-lyric (M2L) generation is an emerging task focuses on creating lyrics that align with the rhythm and structure of a given melody (Sheng et al., 2021). Existing M2L works usually produce lyrics from scratch with little user input Sheng et al. (2021); Tian et al. (2023); Ding et al. (2024), limiting the incorporation of user communication intent. While there are many practical applications revising plain texts to lyrics, under the context of lyrics translation Guo et al. (2022), style transfer (e.g., generating rap lyrics from news articles Nikolov et al. (2020)), and AI-assisted lyric refinement Zhang et al. (2022b). They did not consider melody constraints and thus cannot generate lyrics that are singable with a given melody. A revision framework which edits plain text into melody-aligned lyrics is yet to be explored.
Furthermore, state-of-the-art LLMs et al. (2023) struggle with producing singable lyrics that align precisely with a specific melody. For instance, as illustrated in Figure 1, GPT-4 generates coherent lyrics but fails to synchronize syllables with the last four music notes. Moreover, important words like ‘arm’ and ‘eye’ are paired with less significant notes, disrupting the overall musical flow and resulting in low prosody. On the other hand, prior works on M2L generation Zhang et al. (2022a); Ram et al. (2021) either perform subpar on lyric quality, or/and struggle to generate full-structured lyrics, such as verses or choruses, with good singability. For example, Sheng et al. (2021) employed masked sequence-to-sequence pre-training and attention-based alignment modeling but overlooked prosody and singability, which limited lyrical quality. While recent research (Tian et al., 2023) has enhanced singability by aligning stressed syllables with prolonged musical notes, it still fails to consider prosody, which emphasizes the alignment of key lyrical words and prominent musical notes – a crucial aspect for lyrics-melody alignment.
Addressing these challenges, we propose a novel revision framework, Reffly, which transfers a draft prose to structured and singable lyrics align with a piece of melody. To enhance melody-lyric alignment, we develop a training-free heuristic for capturing prominent lyrical words and musical notes (§$\S$3.1). Since the melody-aligned lyric data is scarce due to copyright constraints, we design an instruction-based mechanism to guide LLMs towards highly singable lyrics by training on a synthetic dataset (§$\S$ 3). Reffly can generate full-length songs with lyrical refrains of verse (develops the song’s plot and message) and chorus (the repeated, memorable musical motif).
@@
  • We propose the first melody-constrained lyric revision framework that, given a predefined melody, transfers an arbitrary text (also referred to as a draft or unsingable lyrics) to a full-length, melody-aligned lyrics with high singability and prosody (also referred to as revised or singable lyrics).
  • We introduce a training-free heuristic for capturing melody-lyrics alignment, semantically and musically, to improve both singability and prosody. Correspondingly, we also contribute a expert labeled dataset with fine-grained annotations of music sheets.  [^1] Anonymous dataset source: https://bit.ly/3X6nCquhttps://bit.ly/3RytMfR
  • In comprehensive experiments across two settings: 1) generation of lyrics from user-specified inputs, and 2) translation of lyrics from Chinese to English, Reffly significantly enhances lyrics-melody alignment and text quality of the generated lyrics, resulting in a 25% and 34% improvements over strong baselines in terms of musicality and overall preference, respectively. [^2] Anonymous demo: https://bit.ly/4fGKWT3
@@ 2.1 What Makes a Good Lyric?
  • Singability is what makes a song easier to sing. For example, it is considered not singable when one single music note maps to a multi-syllable word (e.g., beau-ti-ful) in the lyrics.
@@ 3 Revision Framework for Lyrics
Figure 2 illustrates the inference process of Reffly. To manage the complexity of lyric revision, the revision process is conducted at the sentence level. We iteratively revise each sentence from the unsignable draft to lyric that fits the melody, aligning the prominent words with prominent notes while maintaining the overall coherence.
@@
In this section, we detail each component of Reffly. We developed a training-free heuristic for capturing prominent lyrical words and musical notes (Section §$\S$3.1). Then, §$\S$3.2 introduces revision module for refining unsingable drafts based on musical constraints. Next, §$\S$3.3 provides an overview of the inference process to achieve optimal lyric-melody alignment. Lastly, §$\S$3.4 describes our methods for controlling song structure and generating full-length songs.
3.1 Prominent Musical Notes and Words
Experienced human singers naturally stress prominent words for expressiveness, which align with prominent musical notes Robinson (2005). We introduce the heuristic for identifying prominent notes and words, which are used in generating singable lyric (§$\S$3.3) and the construction of synthetic training data (§$\S$4.1.1).
Prominent Musical Notes
@@ 3.1 Prominent Musical Notes and Words
Inspired by prior research in music theory Caroline Palmer (2006), we developed a heuristic to identify prominent musical notes that stand out in melodies. Our heuristic identifies prominent notes based on three fundamental characteristics of music: Time signature, Rhythm, and Pitch. See Appendix E.2 for details.
Prominent words
Nouns, verbs, adjectives are crucial for effectively conveying meaning through emphasis and recognition (Reikofski, 2015). We validated this concept within the context of lyrics using our constructed datasets, which reveal that nouns, verbs, and adjectives are the most significant parts of speech in terms of prominence: 91.8% of important words are either nouns, verbs, or adjectives. Hence, we identify these part-of-speeches as prominent words (details in Appendix E.3).
3.2 Revision Module
@@
To achieve high-quality lyrics revision, we fine-tuned a revision module with an instruction-based mechanism, as shown in Figure 2, We selected LlaMA2-13b-chat Touvron et al. (2023) as our base model. To effectively transform an unsingable draft into lyrics that fit a given melody while preserving the original meaning, we must address three main tasks:
(1) Syllable planning, which involves generating sentences with the required number of syllables for singability. This task is challenging for LLMs due to difficulties with numerical planning Sun et al. (2023) and the lack of a specific fine-tuning dataset. Moreover, aligning prominent words with prominent notes requires defining which notes and words are prominent.
@@
(2) Aligning prominent words with prominent notes to enhance prosody by ensuring that all prominent words match prominent notes. We propose a "pseudo music constraint" and an instruction template (as illustrate in Figure 3) to improve the syllable planning and word-note matching. The pseudo constraint, derived from generated lyrics, indicates prominent note positions and syllable counts. During training, the model follows pseudo constraints, while in inference, it applies melody constraints. (refer §$\S$§D.2 for more details)
(3) Maintaining coherence for fluent, logical lyrics. Our revision approach operates at the sentence level. To ensure coherence, we provide the model with previously generated lyrics and the song’s title for each training data point.
3.3 Generate Singable Lyric
To generate a full-length singable lyric, we use Reffly to iteratively revise each unsingable sentence. These revised sentences are then seamlessly merged to form the final output.
@@ 3.3 Generate Singable Lyric
As illustrated in figure 2, Reffly takes the melody and each sentence from unsingable draft as input, to produce singable lyrics as the output, with three steps as following.
@@
We extract features from input for further processing. We use our proposed heuristic (refer to §$\S$3.1) to identify prominent musical notes from the melody, encoding them into a melody constraint. We then assemble a prompt using the constraint, title, context, and lyrics draft as input features.
@@
The prompt includes instructions regarding the syllable count to match the melody’s rhythm, steps for refining lyrics, consistency checks with previous context to ensure coherence, and song structure guidance to maintain desired song formats.
@@
To enhance the model capability of generating singable and prominence-aligned lyrics, we break down the process into two sub-steps, and iteratively generate the lyric.
Step 3.1 Revise: We use diverse beam search to generate multiple candidate revisions of the unsingable draft, evaluating each for singability. A lyric is singable if: 1) each note corresponds to one or zero syllables; 2) each syllable in multi-syllable words matches a note no longer than a half-note; 3) multi-syllable words do not cross rests. If no candidates meet these criteria, we restart the process with a rephrased draft.
@@
Algorithm 1 Candidates Selection
@@
#$\vspace{-2mm}\#\text{Ties}=\#\text{Notes}(m)-\#\text{Syllables}(c)\vspace{-2mm}$ (1)
@@
Finally, we select the candidate whose most important words align with prominent notes. If multiple candidates align perfectly, we choose the one most similar to the original sentence based on BERTScore. Zhang et al. (2019).
3.4 Incorporating Full-Song Structure
To generate full-length songs, our framework integrates structure-awareness. During the fine-tuning phase, we get song structure (as introduced in section 2.1) for each lyric line from data source, enabling the model to identify distinctive features of song parts like verses and choruses. As shown in Figure 3, structure tags are embedded directly within the instructions component of our training sequences. Then, as illustrated in Figure 2, during inference (Step 2) , the target structure tags are integrated into the prompt. This strategic incorporation ensures that the generated lyrics align with the melody and are cohesively structured.
@@ 4 Experiments setup
4.1 Dataset Setup
We prepared two datasets: 1) a synthetic training dataset, constructed using our devised strategy to fine-tune the model, and 2) a validation dataset, employed to assess our heuristic (§$\S$ 3.1).
4.1.1 Training dataset
@@
4.1.2 Validation dataset
To evaluate the accuracy of our prominent note extraction heuristic, we created a dataset consisting of 100 song clips annotated by professional musicians, highlighting all the important notes in each melody. This dataset encompasses a diverse range of music styles, including Jazz, Country, Blues, Folk, Pop, and Comedy. An exemplary data point is shown in Figure 6 at appendix.
@@ 4.2 Tasks Setup
4.3 Baselines
We compare our frameworks performance with following models: 1. Lyra is an unsupervised, hierarchical melody-conditioned lyric generator that can generate high-quality lyrics with content control without training on melody-lyric data Tian et al. (2023). 2. SongMass is an LLM design that leveraging masked sequence to sequence (MASS) pre-training and attention based alignment modeling for lyric-to-melody and melody-to-lyric generation Sheng et al. (2021). 3. GPT-4 is a strong versatile LLM et al. (2023) to compare with. We utilize few-shot prompt to provide a template and instruct the model to follow it. 4. Reffly-S represents a variant of our proposed model, from which we have removed the candidate selection algorithm ( introduced in section 3 ). 5. Reffly-I is a variation of our proposed model, where we exclude the instruction component during training(§$\S$ 4.1.1 ). [^6] We describe details of baselines in Appendix C.1.
4.4 Automatic Evaluation Setup
@@
4.5 Human Evaluation Setup
Qualification Task is conducted to select qualified annotators with 1) adequate expertise in song and lyric annotation, and 2) reliable focus on the Mechanical Turk platform. Additional details on the qualification task are provided in Appendix B.1.
Annotation Task Our annotation process is comparative, with annotators reviewing groups of songs produced by various systems that share the same melody and title. All baseline models were assessed. At least three workers annotated each song, rating the lyrics’ quality on a 1-5 Likert scale across five categories. For musicality, the workers assessed prosody (whether prominent words were exaggerated by melody), intelligibility (whether the lyric content was easy to understand when listen to it), and singability (how clearly the lyrics could be understood). In terms of text quality, they evaluated coherence and creativity. The average inter-annotator agreement in terms of Pearson correlation was 0.69.
@@ 4.5 Human Evaluation Setup
Table 1: Evaluation Results for the Arbitrary Generation task. Reffly and its variants (Reffly-S and Reffly-I) consistently outperform other models across most metrics, both in automatic and human evaluations.

### v2

@@ Abstract
Automatic melody-to-lyric (M2L) generation aims to create lyrics that align with a given melody. While most previous approaches generate lyrics from scratch, revision—editing plain text draft to fit it into the melody—offers a much more flexible and practical alternative. This enables broad applications, such as generating lyrics from flexible inputs (keywords, themes, or full text that needs refining to be singable), song translation (preserving meaning across languages while keeping the melody intact), or style transfer (adapting lyrics to different genres). This paper introduces Reffly (REvision Framework For LYrics), the first revision framework for editing and generating melody-aligned lyrics. We train the lyric revision module using our curated synthesized melody-aligned lyrics dataset, enabling it to transform plain text into lyrics that align with a given melody. To further enhance the revision ability, we propose training-free heuristics aimed at preserving both semantic meaning and musical consistency throughout the editing process. Experimental results demonstrate the effectiveness of Reffly across various tasks (e.g. lyrics generation, song translation), showing that our model outperforms strong baselines, including Lyra Tian et al. (2023) and GPT-4, by 25% in both musicality and text quality.
@@ 1 Introduction
Music acts as an important universal language that facilitates social connection and strengthens community involvement Cross (2009). Automatic melody-to-lyric (M2L), creating lyrics that are aligned with a given melody, has emerged as a promising task and received interest by the AI community, because it makes the process of music creation more accessible to a wider audience.(Sheng et al., 2021; Tian et al., 2023; Ding et al., 2024).
In practice, amateur songwriters may wish to craft lyrics that goes with their favorite melody with desired content Tian et al. (2023); Qian et al. (2023), or translate songs into different languages for a wider audience or adapt existing lyrics to fit a different melody Longshen et al. (2023); Nikolov et al. (2020). However, existing M2L approaches and AI-assisted songwriting frameworks fall short to support these use cases, due to insufficient control over sentence-level semantic. Most prior works generate lyrics with zero or limited user input such as keywords or topics (Sheng et al., 2021; Tian et al., 2023; Ding et al., 2024). Some prior works rely on in-filling text templates without providing sufficient automacy to the user (Zhang et al., 2022c; Liu et al., 2020). In contrast, our revision framework refining plain text into melody-aligned lyrics offers greater flexibility and control.
In addition, both state-of-the-art LLMs and prior works on lyrics generation struggle with producing singable lyrics that align well with a specific melody. For instance, as illustrated in Figure 1, ChatGPT-4 OpenAI (2023) generates coherent lyrics but fails to synchronize syllables with the last four music notes. Moreover, prominent words like ‘arm’ and ‘eye’ are paired with less prominent notes, disrupting the overall musical flow and resulting in low prosody. Similarly, prior works on lyrics generation generation either don’t consider melody as a constraint Zhang et al. (2022b); Ram et al. (2021); Zhang et al. (2022c); Liu et al. (2020); Sun et al. (2022), or overlooked the important relationship between prominent note in melody and prominent lyric words (Sheng et al., 2021; Tian et al., 2023; Qian et al., 2023), lowering the generated lyric quality.
Addressing these challenges, we propose a novel revision framework, Reffly, which transfers a draft prose to structured and singable lyrics align with a piece of melody. To enhance melody-lyric alignment, we develop a training-free heuristic for capturing prominent lyrical words and musical notes (§$\S$3.1). Since the melody-aligned lyric data is scarce due to copyright constraints, we design an instruction-based mechanism to guide LLMs towards highly singable lyrics by training on a synthetic dataset (§$\S$ 4.1). Reffly can generate full-length songs with lyrical verses that develop the song’s plot and message, and choruses that repeat a memorable musical motif.
@@
  • We propose the first melody-constrained lyric revision framework that, given a predefined melody, transfers an arbitrary text (also referred to as a draft or unsingable lyrics) to a full-length, melody-aligned lyrics with high singability and prosody (also referred to as revised or singable lyrics), with sentence-level semantic control.
  • We introduce a training-free heuristic for capturing melody-lyrics alignment, semantically and musically, to improve both singability and prosody. Correspondingly, we also contribute a expert labeled dataset with fine-grained annotations of music sheets.  [^1] Dataset source: https://bit.ly/3X6nCquhttps://bit.ly/3RytMfR
  • In comprehensive experiments across two settings: 1) generation of lyrics from user-specified inputs, and 2) translation of lyrics from Chinese to English, Reffly significantly enhances lyrics-melody alignment and text quality of the generated lyrics, resulting in a 25% and 34% improvements over strong baselines in terms of musicality and overall preference, respectively. [^2] Demo: https://bit.ly/4fGKWT3
@@ 2.1 What Makes a Good Lyric?
  • Singability is what makes a song easier to sing. For example, it is considered not singable when one single music note maps to a multi-syllable word (e.g., beau-ti-ful) in the lyrics Tian et al. (2023).
@@ 3 Revision Framework for Lyrics
Figure 2 illustrates the inference process of Reffly. To manage the complexity of lyric revision, the revision process is conducted at the sentence level. We iteratively revise each sentence from the unsingable draft to lyric that fits the melody, aligning the prominent words with prominent notes while maintaining the overall coherence.
@@
In this section, we detail each component of Reffly. We develop a training-free heuristic for capturing prominent lyrical words and musical notes (Section §$\S$3.1). Then, §$\S$3.2 introduces our lyrics revision module that refines unsingable drafts based on musical constraints. Last, §$\S$3.3 provides an overview of the inference process to achieve optimal lyric-melody alignment.
3.1 Aligning Melody with Lyrics
Building on the way experienced singers emphasize certain lyrics to enhance their connection with the melody for musical expressiveness Robinson (2005), we develop a heuristic to align prominent words with prominent musical notes. [^3] It is not feasible to use a neural network-based method for aligning prominent words with prominent musical notes due to the lack of such annotated data. This subsection outlines the process of identifying these prominent notes and words, for which we constructed an expert-annotated dataset to evaluate their effectiveness (see §$\S$5.1 for results). These heuristics are then used in lyrics generation (§$\S$3.3) and the construction of synthetic training data (§$\S$4.1).
Extracting Prominent Musical Notes
@@ 3.1 Aligning Melody with Lyrics
We identify prominent musical notes that stand out in melodies based on three fundamental characteristics of music: Time signature, Rhythm, and Pitch Caroline Palmer (2006). A musical note is considered prominent if it appears on a stressed beat location as defined by its time signature, exhibits syncopation, or having a large pitch jump from the proceeding notes (More details in Appendix E.2).
Extracting Prominent words from Lyrics
According to Reikofski (2015), nouns, verbs, adjectives are crucial for effectively conveying meaning. Therefore, we identify nouns, verbs, and adjectives that are non-stop words as prominent words.
Assessing the Accuracy
To the best of our knowledge, we are the first to apply computational algorithms to identify prominent notes and words. Therefore, to evaluate the accuracy of our heuristic, we collected a validation dataset consisting of 100 song clips, each containing three to five musical phrases, annotated by professional musicians marking all the important notes in each melody. This dataset covers a diverse range of music styles, including Jazz, Country, Blues, Folk, Pop, and Comedy. Figure 6 shows an exemplary data point.
3.2 Lyrics Revision Module
@@
To achieve high-quality lyrics revision, we fine-tuned a LlaMA2-13b-chat model Touvron et al. (2023) to effectively transform an unsingable draft into lyrics that fit a given melody while preserving the original meaning. We address three main tasks:
(1) Syllable planning: Generating sentences with the necessary syllable count to ensure singability. (2) Aligning prominent words with notes: Matching the stressed syllable of prominent words with prominent notes to enhance prosody. (3) Maintaining local and global coherence: Ensuring smooth transitions between sentences for local coherence and capturing global structures, such as verse and chorus, for structure-awareness. Addressing these tasks is challenging for LLMs, which are known to struggle with numerical planning Sun et al. (2023). Furthermore, we face a lack of labeled datasets to train a supervised model that learned to map the prominence in lyrics and melody.
@@
To empower the model with the first two abilities, we propose a “pseudo music constraint” (blue box of Figure 3) to improve the syllable planning and word-note matching. The pseudo constraint, derived from generated lyrics, indicates prominent note positions and syllable counts. During training, the model follows pseudo constraints, while in inference, it applies melody constraints. (refer to §$\S$§D.2 for more details).
To maintain both local and global coherence, we introduce an instruction template (purple box of Figure 3). To ensure local, sentence-level, coherency, we provide the model with previously generated lyrics and the song’s title as context for each training data point.To enable full-length song generation, our framework incorporates structure-awareness by embedding song structure into the fine-tuning phase. During this phase, we use song structure information ( introduced in section 2.1) from the data source for each lyric line, allowing the model to recognize features of different song structures like verses and choruses. Structure tags are embedded within the instruction component and integrated into the prompt.
3.3 Generate Lyrics at Inference Time
@@
As is illustrated in figure 2, Reffly takes the melody and each sentence from unsingable draft as input, to produce singable lyrics as the output, with three steps as following.
@@
[…]
@@
We assemble instructions that specify various features for the desired lyrics, such as matching the syllable count given melody, refining lyrics, providing previous context to ensure coherence, and guidance to maintain desired song structures. An example input prompt can be found in  Figure 9 in Appendix.
@@
To enhance the model’s ability to generating singable, prominence-aligned lyrics, we break down the process into two sub-steps, iteratively generating the lyrics.
Step 3.1 Revise: We adopt diverse beam search Vijayakumar et al. (2016) to generate multiple candidate revisions of the unsingable draft, evaluating each for singability. A lyric is singable if: 1) each note corresponds to one or zero syllables; 2) each syllable in multi-syllable words matches a note no longer than a half-note; 3) multi-syllable words do not cross rests. If no candidates meet these criteria, we restart the process with a rephrased draft.
@@
Algorithm 1 Candidate Selection
@@
#$\vspace{-2mm}\#\text{Ties}=\#\text{Notes}(m)-\#\text{Syllables}(c)$ (1)
@@
Finally, we select the candidate whose most important words align with prominent notes. If multiple candidates align perfectly, we choose the one most similar to the original sentence based on BERTScore Zhang et al. (2019).
@@ 4 Experiments setup
4.1 Synthetic Training Dataset
@@ 4.2 Tasks Setup
4.3 Compared Models
We compare our framework with two baselines and introduce two ablation variations of Reffly to validate each component.
Baselines. 1. Lyra is an unsupervised, hierarchical melody-conditioned lyric generator that can generate high-quality lyrics with content control without training on melody-lyric data Tian et al. (2023). 2. SongMass is an LLM design that leveraging masked sequence to sequence (MASS) pre-training and attention based alignment modeling for lyric-to-melody and melody-to-lyric generation Sheng et al. (2021). 3. GPT-4 is a strong versatile LLM OpenAI (2023) to compare with. We utilize few-shot prompt to provide a template and instruct the model to follow it.
Variations. 4. Reffly w/o S. is a variant of our proposed framework without the candidate selection algorithm (shown in the green and yellow boxes of Step 3.2 in Figure 2). 5. Reffly w/o I. excludes the instruction component during training (purple and red boxes in Figure 3 ). [^6] We describe details of baselines in Appendix C.1.
4.4 Evaluation Setup
We conduct both automatic and human evaluations to assess our framework. While human evaluation is more reliable, it is difficult to scale and reproduce. Therefore, we use widely-adopted metrics like diversity, perplexity, and BERTScore to evaluate creativity, smoothness, and semantic similarity between generated lyrics and initial drafts (Sheng et al., 2021; Tian et al., 2023).
4.4.1 Automatic Evaluation
@@
Table 1: Evaluation Results for the Arbitrary Generation task. Reffly and its variants (Reffly w/o S. and Reffly w/o I.) consistently outperform other models across most metrics, both in automatic and human evaluations.
4.4.2 Human Evaluation
Annotation Task
We conducted a qualification task to select annotators with expertise in song and lyric annotation on Mechanical Turk. Additional details on the qualification process are provided in Appendix B.1. Our annotation process is comparative, with annotators reviewing groups of songs produced by various systems that share the same melody and title. All baseline models were assessed. At least three workers annotated each song, rating the lyrics’ quality on a 1-5 Likert scale across five categories. For musicality, the workers assessed prosody (whether prominent words were exaggerated by melody), intelligibility (whether the lyric content was easy to understand when listen to it), and singability (how clearly the lyrics could be understood). In terms of text quality, they evaluated coherence and creativity. Annotators evaluated structural clarity (whether the verse-chorus structure is clear) structural-aware generation and assessed translation quality in song translation. The average inter-annotator agreement in terms of Pearson correlation was 0.69.